Consider the fully developed flow of glycerin at 40 °C through a 70-mlong, 4-cm-diameter, horizontal, circular pipe. If the flow velocity at the
centerline is measured to be 6 m/s, determine the velocity profile and the
pressure difference across this 70-m-long section of the pipe, and the
useful pumping power required to maintain this flow.
1
Expert's answer
2020-09-03T14:10:58-0400

solution:-

given data

density of glycerin ("\rho" ) =1252 kg/m^3

viscosity of glycerin("\mu" )=0.3073 kg/m-s

diameter of pipe (D) =0.04 m

length of pipe (L)=70 m

maximum velocity(umax)=6 m/s

the velocity profile in fully developed laminar flow is


"u(r)=u_{max}(1-\frac{r^2}{R^2})"


"=6(1-\frac{r^2}{(0.02)^2})"


"\fcolorbox{red}{yellow}{$u(r)=6(1-2500r^2)m/s$}"


average velocity

"V_{a}=\frac{u_{max}}{2}=\frac{6}{2}=3m/s"


flow rate

"Q=V_aA"


"Q=3\times\pi\times\frac{(0.04)^2}{4}"

"Q=3.77\times10^{-3}m^3/s"

renolds number can be given


"Re=\frac{\rho V_aD}{\mu}"


"=\frac{1252\times3\times0.04}{0.3073}"


"Re=488.9"


"Re<2300"

so its laminar flow.

therefore friction factor

"f=\frac{64}{Re}=\frac{64}{488.9}=0.1309"


and head loss

"H=f\frac{LV_a^2}{D2g}"


"H=\frac{0.1309\times70\times3^2}{0.04\times2\times9.8}"

"H=105.1m"

by using the Bernoulli equation pressure difference can be written as


"\Delta P=P_1-P_2 =\rho g(Z_2-Z_1+H)"


pipe is horizontal so Z1 and Z2 will be zero.


"\Delta P=1252\times9.8(0+105.1)"


"\fcolorbox{green}{yellow}{$\Delta P=1291\times10^3 Pa$}"


pumping power can be given as


"pumping\space power = Q \times\Delta P"

"=3.77\times10^{-3}\times1291\times10^{3}"


"\fcolorbox{green}{yellow}{$pumping \space power=4.87\times10^3 W$}"